The length of a rectangular room is 6 feet longer than twice the width.If the room's perimeter is 144 feet, what are the room's dimension's?
Answer by checkley77(12844)

L=2W+6
P=2L+2W
144=2(2W+6)+2W
144=4W+12+2W
144=6W+12
6W=144-12
6W=132
W=132/6
W=22 FT. ANS. FOR THE WIDTH.
L=2*22+6
L=44+6=50 FT. ANS. FOR THE LENGTH.
PROOF:
144=2*50+2*22
144=100+44
144=144